w=2 y=6 x=5
I have two problems
yx2 and (yx)2
are these both the same?
and is the yx 65?
Help!!!!!
Wyatt
i think the name of the math book is mathematics i forgot to bring it home
: w=2 y=6 x=5
I have two problems
yx2 and (yx)2
are these both the same?
and is the yx 65?
Help!!!!!
Wyatt
i think the name of the math book is mathematics i forgot to bring it home

Answer by Alan3354(1449)

w=2 y=6 x=5
I have two problems
yx2 and (yx)2
are these both the same?
and is the yx 65?
-------------
yx2 is y*x*x, = 6*5*5
That's 150
(yx)2 is yx*yx = 6*5*6*5
That's 900
Obviously they're not the same.